Selected operating data for two divisions of Outback Brewing, Ltd., of Australia are given below: Division Queensland New South Wales Sales Average operating assets Net operating income Property, plant, and equipment (net) S 864,000 S 2,016,000 $ 540,000 630,000 B0,640 S 247,000 197,000 $60,480 S Required 1. Compute the rate of return for each division using the return on investment (ROI) formula stated in terms of margin and turnover. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Margin Turnover ROI Queensland division times New South Wales division times 2. Which divisional manager seems to be doing the better job? Queensland division O New South Wales division ReferenceseBook & Resources
Answer #1

Margin =Net operating income/Total Sales

Turnover=total sales/Average Asset

ROI =Margin * Turnover

Quuen Land Division

Margin =60480/864000=7.00%

Turnover=864000/540000=1.6

ROI=7.00%*1.6=11.20%

New South wale Division

Margin =80640/2016000=4.00%

Turnover=2016000/630000=3.2 times

ROI=4%*3.2=12.80%

2

New South wale Division As the ROI is greater